Related Papers

References

[1] G. Cooperman, “TOP-C: Task-Oriented Parallel C for Distributed and Shared Memory”, Workshop on Wide Area Networks and High Performance Computing, Lecture Notes in Control and Information Sciences 249, 1999, Springer Verlag, pp. 109–118.

[2] G. Cooperman, “Practical Task-Oriented Parallelism for Gaussian Elimination in Distributed Memory”, Linear Algebra and its Applications 275-276, 1998, pp. 107–120.

[3] G. Cooperman, “TOP-C: A Task-Oriented Parallel C Interface”, 5th International Symposium on High Performance Distributed Computing (HPDC-5), 1996, IEEE Press, pp. 141–150.

[4] G. Cooperman, “GAP/MPI: Facilitating Parallelism”, Proc. of Second DI-MACS Workshop on Groups and Computation II 28, DIMACS Series in Discrete Mathematics and Theoretical Computer Science, L. Finkelstein and W.M. Kantor (eds.), AMS, Providence, RI, 1997, pp. 69–84.

[5] G. Cooperman, “STAR/MPI: Binding a Parallel Library to Interactive Symbolic Algebra Systems”, Proc. of International Symposium on Symbolic and Algebraic Computation (ISSAC ’95), ACM Press, 1995, pp. 126–132.

[6] G. Cooperman, L. Finkelstein, M. Tselman and B. York, “Constructing Permutation Representations for Matrix Groups”, J. Symbolic Computation 24, 1997, pp. 1–18.

[7] G. Cooperman and G. Havas, Practical parallel coset enumeration, Proc. of Workshop on High Performance Computation and Gigabit Local Area Networks, G. Cooperman, G. Michler and H. Vinck (eds.), Lecture notes in control and information sciences 226, Springer Verlag, pp. 15–27.

[8] G. Cooperman, G. Hiss, K. Lux, and Jürgen Müller, The Brauer tree of the principal 19-block of the sporadic simple Thompson group, J. of Experimental Mathematics 6(4), 1997, pp. 293–300.

[9] G. Cooperman and M. Tselman, New Sequential and Parallel Algorithms for Generating High Dimension Hecke Algebras using the Condensation Technique, Proc. of International Symposium on Symbolic and Algebraic Computation (ISSAC ’96), ACM Press, 155–160.